### v2

@@ Abstract
Despite their proficiency in math tasks, the mechanisms underlying LLMs’ mathematical reasoning abilities remain a subject of debate. Recent studies suggest that chain-of-thought (CoT) prompts can bolster mathematical reasoning by encouraging LLMs to employ human-like logical reasoning (System 2), enabling them to excel on the Cognitive Reflection Test (CRT). To assess whether LLMs genuinely possess System 2-like logical reasoning, we introduced targeted modifications to CRT problems. Our findings reveal that, despite the use of CoT prompts, mainstream LLMs, including the latest o1-preview model, continue to exhibit a significant error rate. Further analysis indicates that they predominantly rely on System 1-like intuitive reasoning and pattern matching derived from training data, rather than demonstrating mastery of mathematical thinking. This discovery challenges the prevailing notion that LLMs possess genuine logical reasoning abilities and that CoT can enhance them. Consequently, this work may temper overly optimistic projections regarding LLMs’ advancement toward artificial general intelligence.
@@ 1 Introduction
Large language models (LLMs) are hailed as the dawn of Artificial General Intelligence (AGI)[1]. LLMs such as ChatGPT[2], GPT-4[1], Claude[3], Gemini[4], GLM[5], and o1-preview[6] have garnered significant attention from both academia and industry, have demonstrated tremendous potential in various fields such as education[7], healthcare[8, 9], coding[10], and social governance[11]. One of the key reasons for this is the ‘emergence phenomenon’[12], enabled through massive training data and vast parameter counts: LLMs exhibit unexpected competence in tasks they were not specifically trained for.
Using mathematical skills as an illustration, not all LLMs have received specific training for solving complex mathematical problems during their development. Nonetheless, most of them have shown remarkable abilities to tackle these issues. Moreover, by using the CoT method, the ability of LLMs to solve mathematical problems can be further enhanced [13, 14, 15, 16, 17]. However, due to the interpretability challenges posed by large-scale neural networks, there is no scientific consensus yet on the origins of the mathematical capabilities of LLMs. Meanwhile, recent research has questioned the effectiveness of the CoT method, suggesting that CoT provides strong performance benefits primarily on tasks involving math or logic, with much smaller gains on other types of tasks[18].
@@
Specifically, Thilo Hagendorff et al. [19] have demonstrated through experiments that the CoT method can effectively help LLMs handle the pitfalls in Cognitive Reflection Test (CRT) problems[20]. CRT problems are some well-crafted math or logic problems that human testers often get wrong due to intuitive thinking (System 1[21, 22, 23, 24]). By using the CoT method, LLMs are prompted to rely more on human-like logical reasoning (System 2[21, 22, 23, 24]), enhancing their accuracy in solving these problems. Advanced models like GPT-4 have even achieved higher accuracy than humans in these CRT tasks. However, the authors also raised speculative questions in their work, suggesting, “It is possible that some models encountered enough examples in their training to solve them ‘from memory’”[19]. Our study aims to investigate this speculation through empirical research.
@@
We repeated and improved upon the experiment conducted by the previous study [19], making targeted (yet as simple as possible) modifications to the original CRT problems. If LLMs genuinely possess the intrinsic capability to comprehend mathematical logic, as is widely hypothesized, their accuracy in responding to the modified problems should not experience a marked decrement. However, our experiments revealed a distinctly opposing result: even when employing the CoT approach, prominent LLMs, including the latest o1 model, continued to manifest a considerable error rate for the modified problems. Further analysis indicates that LLMs may have not developed a logic reasoning proficiency akin to System 2, nor have they acquired comprehensive mathematical cognitive abilities, including principles of mathematical operations and algorithm selection. They predominantly resort to a methodology reminiscent of System 1, which involves matching and producing responses based on the similarity between user inquiries and training data during the process of text generation. This investigation may serve to temper the overly optimistic anticipations regarding the effectiveness of CoT and the competencies of LLMs in approximating AGI.
@@ 2 Experiment I: Changing the numbers in a problem without altering its description and principle
We first replicated the experiment on the CRT3 dataset from Thilo Hagendorff’s study[19], which comprises 50 mathematical test problems. We introduced three types of modifications to the test problems, as delineated in Table 1. Type A represents the original problems in the CRT3 test, characterized as typical exponential growth problems, and incorporates three numbers in the problem statement. Type B modifies two of these numbers, while type C modifies all three. Type D replaces two numbers with letters, transforming the arithmetic problem into an algebraic one. To prevent LLMs from being unsure how to handle algebraic symbols, we appended a suffix for guidance: “X and Y are both numbers, you can use them to represent the final answer.” The modifications in Types B, C, and D do not alter the description of the original problem (Type A). Thus, the underlying mathematical principle of the original problem remains unchanged. To uncover the problem-solving strategies of LLMs, we appended explicit instructions (’Please think step by step.’) to the end of each of the four types of problems, encouraging the use of the CoT method.
@@
D Algebraic Transformation In a city, a virus is spreading, causing the total number of infected individuals to triple every 3X days. If it takes 6Y days for the entire city’s population to be infected, how many days would it require for 1/27 of the people to become infected? X and Y are both numbers, you can use them to represent the final answer. Please think step by step.
@@
We evaluated five mainstream open-source and commercial models, manually analyzing and recording their accuracy in responding to the problems. A response was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices A ). To ensure the reliability of our experimental results, we conducted three replicates of the experiments and computed the average for each data set. The results are shown in Figure 1. In the case of Type A problems, the highest accuracy reached 100.00%, the lowest was 54.00%, and the average accuracy was 86.80%. For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.00%, the lowest 25.33%, and the average 68.53%($\delta=18.27\%;\chi^{2}(1)=23.00;P<0.001$). For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.00%, the lowest at 10.66%, and the average 53.06%($\delta=15.47\%;\chi^{2}(1)=11.91;P<0.001$). The accuracy dropped sharply once again when the arithmetic problems were transformed into algebraic ones (Type D). The best-performing generic model, GPT-4, achieved an accuracy of only 29.33%, whereas ChatGPT 3.5 scarcely provided correct responses to any problems. The other models only got around ten problems right, reducing the average accuracy down to 20.93%($\delta=32.13\%;\chi^{2}(1)=54.00;P<0.001$).
Result Analysis of Experiment I: The modifications in types B, C, and D only altered the specific numbers in the original problems without changing the problems’ mathematical principles and computational rules. For a computer program, merely changing the input parameters without modifying the computation process should not lead to a decrease in accuracy. However, the performance of LLMs differs significantly from that of computer programs. We analyzed all incorrect responses provided by the five LLMs for each type of problem and found that in type B, errors arising from mathematical operations (e.g., 16/2=4) constituted 6.78%, while those due to errors in calculation steps (e.g., omitting a step or altering the original calculation method) accounted for 83.47%, and errors caused by both factors accounted for 9.75%. In type C, errors solely due to mathematical operations accounted for 5.11%, errors due to calculation steps accounted for 85.51%, and errors caused by both factors accounted for 9.38%. In type D, errors due to mathematical operations accounted for 2.19%, errors due to calculation steps accounted for 95.45%, and errors caused by both factors accounted for 2.36%.
@@ 3 Experiment II: Modifying the principle of a problem while maintaining similarity in its description
In this experiment, we substantially altered the fundamental principles of the mathematical problems, endeavoring to preserve descriptions that closely mirrored the original versions. Subsequently, we investigated whether LLMs persisted in utilizing their prior problem-solving strategies or adapted their methods to the modified problem contexts throughout the problem-solving process. We conducted three distinct sub-experiments corresponding to the CRT1, CRT2, and CRT3 datasets.
@@ 3.1 For CRT1: Transforming Additively Separable Problems into Non-Additively Separable Problems
For instance, Figure 2 illustrates the crucial distinction between an original CRT1 problem and its modified version. In the original problem, the total cost is simply the sum of the prices of two items. However, in the modified version, reaching the two items from the starting point through a shared segment must be deducted from the overall distance calculation. We use the string similarity calculation function within the difflib module of Python to quantify the degree of similarity in textual expression between the modified problem and the original one, resulting in an average similarity of 75.91%. Despite the similar wording in the problem statements, the underlying mathematical principles are fundamentally distinct. To reveal how LLM approaches these problems, we included a CoT prompt for both the original and the modified problems.
@@
We evaluated five mainstream open-source and commercial LLMs and analyzed their accuracy in answering both the original problems and the modified problems. An answer was deemed correct only if both the reasoning steps and the conclusion were accurate. For each problem, we replicated the experiment three times, averaging the results to ensure robustness. Figure 3 illustrates our findings. For the original CRT1 problems, four LLMs achieved a perfect accuracy of 100%, with GLM-4 closely following at 95.33%, yielding an overall accuracy of 99.07%. However, when confronted with the modified problems, the accuracy significantly declined. GLM-4 recorded the highest accuracy rate, albeit a mere 5.33%. Claude3 and Gemini-1.5 failed to provide correct responses to any of the modified problems. On average, the five LLMs attained only 1.47% accuracy($\delta=97.60\%;\chi^{2}(1)=472.41;P<0.001$). By analyzing incorrect answers, we found that 95.30% of errors were attributed to the application of the original problem’s solution method to the modified problem, whereas 4.70% resulted from incorrect application of a new method or computational errors.
Figure 3: Accuracy of the five LLMs when answering the original and the modified CRT1 problems.
@@
As illustrated in Figure 4, we take the first problem in the original CRT2 dataset as an example. In the original problem, all workers work together, and the total amount of product is related to the number of workers. In the modified problem, working together is changed to walking together, therefore, the total distance walked is independent of the number of people. We calculated the average similarity using the same function as in the previous sub-experiment, yielding 83.67%. Although the descriptions of the original problem and the modified one are similar, the underlying mathematical principles are entirely different. To uncover the problem-solving strategies of LLMs, we added CoT prompts to both the original and the modified problems.
@@
Figure 5: Accuracy of the five LLMs when answering the original and the modified CRT2 problems.
As illustrated in Figure 5, the experimental findings demonstrate a significant decrease in the accuracy of various models. Claude3’s accuracy plummeted from an initial level of 62.00% to 22.00% ($\delta=40.00\%;\chi^{2}(1)=14.82;P<0.001$). Similarly, GPT-4 experienced a dramatic reduction from 94.00% to 33.33%($\delta=60.67\%;\chi^{2}(1)=37.20;P<0.001$). Gemini-1.5 experienced the most substantial decline, dropping from 94.67% to 19.33%($\delta=75.34\%;\chi^{2}(1)=54.85;P<0.001$). GLM-4 also encountered a considerable decrease, with its accuracy falling from 65.33% to 28.00%($\delta=37.33\%;\chi^{2}(1)=12.54;P<0.001$). ChatGPT 3.5 exhibited a smaller change, declining from 50.67% to 34.67%($\delta=16.00\%;\chi^{2}(1)=2.00;P=0.157$).
Upon scrutinizing all erroneous answers to the modified problems, it was observed that 52.81% of errors employed the original CRT2 methodology to address the modified problems. This finding suggests that the LLMs perceive the modified problems as akin to the originals due to their descriptive similarity, consequently relying on the original solution strategy.
@@ 3.3 For CRT3: Transforming Exponential Growth Problems into Linear Growth Problems
Figure 6 presents the first problem in the CRT3 dataset as an illustrative example. In the original problem, the number of viruses doubles each day, resulting in exponential growth of the total number. In contrast, the modified problem features a constant daily increase in the number of viruses, which leads to linear growth. We calculated the average similarity using the same function, yielding a result of 88.89%. To further elucidate the problem-solving strategies of the LLM, we added a CoT prompt for both the original and the modified problem.
@@
The experimental results in Figure 7 demonstrate that Gemini-1.5 achieved the highest accuracy of 100% for the original CRT3 problems. In contrast, ChatGPT 3.5 recorded the lowest accuracy at 54%, with an average accuracy of 86.80%. Following the modification of the problems, the average accuracy decreased significantly to 12.53% ($\delta=74.27\%;\chi^{2}(1)=272.83;P<0.001$). The accuracy rate of Gemini-1.5 decreased to 20.00%($\delta=80.00\%;\chi^{2}(1)=63.38;P<0.001$), while Claude3’s accuracy was merely 1.33% ($\delta=92.00\%;\chi^{2}(1)=81.23;P<0.001$).
The principal reason for the decline in the average accuracy is that these models continued to employ the problem-solving approach of the original problem when confronted with the modified problem. The proportions of Claude3, ChatGPT 3.5, GPT-4, Gemini-1.5, and GLM-4 that identified the modified problem as one of exponential growth were 81.69%, 77.90%, 85.88%, 72.49%, and 71.49%, respectively, although it should have been identified as a linear growth problem. Among these, the proportion of problems that were solved entirely using the original method—yielding the same answer as the original problem—constituted 37.85%, 31.83%, 82.70%, 55.73%, and 52.17% of all incorrect answers, with an average of 52.06%. This result shows that more than half of the errors were due to using the original method.
Figure 7: Accuracy of the five LLMs when answering the original and the modified CRT3 problems.
Result Analysis of Experiment II: The modifications to CRT1, CRT2, and CRT3 altered the mathematical principles underlying these problems with a effort to preserve the similarity of their statements. The experimental findings reveal that LLMs continue to rely on their inherent problem-solving methodologies, even when provided with CoT prompts as guidance. This observation further substantiates our hypothesis that generative LLMs operate similarly to a word chain game; they can address some mathematical issues probably because their training datasets include analogous problems. Consequently, LLMs tend to adopt problem-solving strategies based on the similarity of problem statements rather than a genuine understanding of the mathematical principles involved.

### v4

@@ Abstract
The cognitive mechanism by which Large Language Models (LLMs) solve mathematical problems remains a widely debated and unresolved issue. Currently, there is little interpretable experimental evidence that connects LLMs’ problem-solving with human cognitive psychology. To determine if LLMs possess human-like mathematical reasoning, we modified the problems used in the human Cognitive Reflection Test (CRT). Our results show that, even with the use of Chains of Thought (CoT) prompts, mainstream LLMs, including the latest o1 model (noted for its reasoning capabilities), have a high error rate when solving these modified CRT problems. Specifically, the average accuracy rate dropped by up to 50% compared to the original questions. Further analysis of LLMs’ incorrect answers suggests that they primarily rely on pattern matching from their training data, which aligns more with human intuition (System 1 thinking) rather than with human-like reasoning (System 2 thinking). This finding challenges the belief that LLMs have genuine mathematical reasoning abilities comparable to humans. As a result, this work may adjust overly optimistic views on LLMs’ progress towards artificial general intelligence.
@@ 1 Introduction
Large language models (LLMs) are considered the dawn of Artificial General Intelligence (AGI) [1]. Models such as ChatGPT[2], GPT-4[1], Claude[4], Gemini[3], GLM[5], and o1-preview[23] have garnered considerable attention from both academia and industry. These models exhibit significant potential across various fields, including education[6], healthcare[7, 8], coding[9], and social governance[10]. This is partly attributed to the ‘emergence phenomenon’ [12], which allows LLMs, due to their large training datasets and numerous parameters, to perform tasks they were not specifically trained for.
Using mathematical skills as an example, most LLMs have demonstrated remarkable abilities to tackle these problems. By leveraging the Chain of Thought (CoT) method, the capacity of LLMs to solve mathematical problems can be further augmented[11, 20, 21, 19, 22]. However, due to the interpretability challenges posed by large-scale neural networks, there is still no scientific consensus on the origins and mechanisms underlying the mathematical capabilities of LLMs.
@@
Conducting psychological measurement experiments on LLMs can help enhance the interpretability of research into LLMs’ thinking. Thilo Hagendorff et al.[13] have demonstrated through experiments that the CoT method can effectively help LLMs handle the pitfalls in Cognitive Reflection Test (CRT) problems[14]. CRT problems are some well-crafted math or logic problems that human testers often get wrong due to intuitive thinking (System 1[15, 16, 18, 17]). By using the CoT method, LLMs are prompted to rely more on human-like logical reasoning (System 2[15, 16, 18, 17]), enhancing their accuracy in solving these problems. Advanced models like GPT-4 have even achieved higher accuracy than humans in these CRT tasks. However, the authors also raised speculative questions in their work, suggesting, “It is possible that some models encountered enough examples in their training to solve them ‘from memory’”.
@@
We repeated and improved upon the experiment conducted by the previous study[13], making targeted (yet as simple as possible) modifications to the original CRT problems. If LLMs genuinely possess the intrinsic capability to comprehend mathematical logic, as is widely hypothesized, their accuracy in responding to the modified problems should not experience a marked decrement. However, our experiments revealed a distinctly opposing result: even when employing the CoT approach, prominent LLMs, including the latest o1 model, continued to manifest a considerable error rate for the modified problems. Further analysis of the incorrect answers indicates that LLMs may have not developed a logic reasoning proficiency akin to System 2, nor have they acquired comprehensive mathematical cognitive abilities. They predominantly resort to a methodology reminiscent of System 1, which involves matching and producing responses based on the similarity between user inquiries and training data during the process of text generation. This investigation may serve to temper the overly optimistic anticipations regarding the effectiveness of CoT and the competencies of LLMs in approximating AGI.
@@ 2 Experiment I: Changing the Numbers in A Problem Without Altering Its Description and Principle
We first replicated the experiment on the CRT3 dataset from Thilo Hagendorff’s study[13], which comprises 50 mathematical problems. We introduced three types of modifications to the test problems, as delineated in Table 1. Type A represents the original problems in the CRT3 test, characterized as typical exponential growth problems, and incorporates three numbers in the problem statement. Type B modifies two of these numbers, while type C modifies all three. Type D replaces two numbers with letters, transforming the arithmetic problem into an algebraic one. To prevent LLMs from being unsure how to handle algebraic symbols, we appended a suffix for guidance: “X and Y are both numbers, you can use them to represent the final answer.” The modifications in Types B, C, and D do not alter the description of the original problem (Type A). Thus, the underlying mathematical principle of the original problem remains unchanged. To uncover the problem-solving strategies of LLMs, we appended explicit instructions (‘Please think step by step.’) to the end of each of the four types of problems, encouraging the use of the CoT method.
@@
D Algebraic Transformation In a city, a virus is spreading, causing the total number of infected individuals to triple every 3X days. If it takes 6Y days for the entire city’s population to be infected, how many days would it require for 1/27 of the people to become infected? X and Y are both numbers, and you can use them to represent the final answer. Please think step by step.
@@
We evaluated five mainstream open-source and commercial LLMs by manually analyzing and recording their answers to the problems, ensuring the precision of our result judgments. An answer was considered correct only if both the problem-solving approach and the final answer were accurate (refer to Appendices A ). To ensure the reliability of our experimental results, we conducted three replicates of the experiments and computed the average for each data set.
The results are shown in Fig. 1. In the case of Type A problems, the highest accuracy reached 100%, the lowest was 54.0%, and the average accuracy was 86.8%. For Type B problems, which involved modifications to two numbers, the accuracy significantly decreased, with the highest being 92.0%, the lowest 25.3%, and the average 68.5%($\delta=18.3\%;\chi^{2}(1)=23.00;P<0.001$). For the problems of Type C, where three numbers were modified, the accuracy further declined, with the highest at 80.0%, the lowest at 10.7%, and the average 53.1%($\delta=15.4\%;\chi^{2}(1)=11.91;P<0.001$). The accuracy dropped sharply once again when the arithmetic problems were transformed into algebraic ones (Type D). The best-performing model, GPT-4, achieved an accuracy of only 29.3%, whereas ChatGPT 3.5 scarcely provided correct answers to any problems. The average accuracy was reduced to 20.9%($\delta=32.2\%;\chi^{2}(1)=54.00;P<0.001$).
Result Analysis of Experiment I: The modifications in types B, C, and D only altered the specific numbers in the original problems without changing the problems’ mathematical principles and computational rules. For a human who possesses mathematical logical thinking and is capable of solving problems through formulating equations or programming, merely changing the input numbers without modifying the mathematic principles should not lead to a decrease in accuracy. However, the performance of LLMs differs significantly from that of a rational human.
As shown in Fig. 2, we analyzed all incorrect answers provided by the five LLMs for each type of problem and found that in type B, errors arising from incorrect solution steps (e.g., omitting a step or altering the original calculation method) constituted 93.2%, while those only due to mathematical calculation (e.g., 16/2=4) accounted for 6.8%. In type C, errors due to incorrect solution steps accounted for 94.9%, errors solely due to mathematical calculation accounted for 5.1%. In type D, errors due to incorrect solution steps accounted for 97.8%, and errors caused solely by mathematical calculation accounted for 2.2%.
Figure 2: Analysis of the proportion of incorrect answers in Experiment I
@@ 3 Experiment II: Modifying the Principle of A Problem While Maintaining Similarity in Its Description
To further substantiate the inferences drawn from Experiment I, we designed a reverse experiment. In this experiment, we substantially altered the fundamental principles of the mathematical problems, endeavoring to preserve descriptions that closely mirrored the original versions. Subsequently, we investigated whether LLMs persisted in utilizing their prior problem-solving strategies or adapted their methods to the modified problem contexts throughout the problem-solving process. We conducted three distinct sub-experiments corresponding to the CRT1, CRT2, and CRT3 datasets.
@@ 3.1 For CRT1: Transforming Additively Separable Problems into Non-Additively Separable Problems
For instance, Fig. 3 illustrates the crucial distinction between an original CRT1 problem and its modified version. In the original problem, the total cost is simply the sum of the prices of two items. However, in the modified version, reaching the two items from the starting point through a shared segment must be deducted from the overall distance calculation. We use the string similarity calculation function within the difflib module of Python to quantify the degree of similarity in textual expression between the modified problem and the original one, resulting in an average similarity of 75.91%. Despite the similar wording in the problem statements, the underlying mathematical principles are fundamentally distinct. To reveal how LLM approaches these problems, we included a CoT prompt for both the original and the modified problems.
@@
We evaluated five mainstream open-source and commercial LLMs and analyzed their accuracy in answering both the original problems and the modified problems. An answer was deemed correct only if both the reasoning steps and the conclusion were accurate. For each problem, we replicated the experiment three times, averaging the results to ensure robustness.
Fig. 4A illustrates our findings. For the original CRT1 problems, four LLMs achieved a perfect accuracy of 100%, with GLM-4 closely following at 95.3%, yielding an overall accuracy of 99.1%. However, when confronted with the modified problems, the accuracy significantly declined. GLM-4 recorded the highest accuracy rate, albeit a mere 5.3%. Claude3 and Gemini-1.5 failed to provide correct answers to any of the modified problems. On average, the five LLMs attained only 1.5% accuracy($\delta=97.6\%;\chi^{2}(1)=472.41;P<0.001$).
Figure 4: Accuracy of the LLMs when answering the CRT1 problems (A), and the proportion of incorrect answer types in modified CRT1 problems (B).
By analyzing incorrect answers, as shown in Fig. 4B, we found that 95.9% of errors were attributed to the application of the original problem’s solution method to the modified problem (Maintain the original solution steps), whereas 4.1% resulted from new solution steps but with errors.
@@
As illustrated in Fig. 5, we take the first problem in the original CRT2 dataset as an example. In the original problem, all workers work together, and the total amount of product is related to the number of workers. In the modified problem, working together is changed to walking together. Therefore, the total distance walked is independent of the number of people. We calculated the average similarity using the same function as in the previous sub-experiment, yielding 82.33%. Although the descriptions of the original problem and the modified one are similar, the underlying mathematical principles are entirely different. To uncover the problem-solving strategies of LLMs, we added CoT prompts to both the original and the modified problems.
@@
As illustrated in Fig. 6A, the experimental findings demonstrate a significant decrease in the accuracy of various models. Claude3’s accuracy plummeted from an initial level of 62.0% to 22.0% ($\delta=40.0\%;\chi^{2}(1)=14.82;P<0.001$). Similarly, GPT-4 experienced a dramatic reduction from 94.0% to 33.3%($\delta=60.7\%;\chi^{2}(1)=37.20;P<0.001$). Gemini-1.5 experienced the most substantial decline, dropping from 94.7% to 19.3%($\delta=75.4\%;\chi^{2}(1)=54.85;P<0.001$). GLM-4 also encountered a considerable decrease, with its accuracy falling from 65.3% to 28.0%($\delta=37.33\%;\chi^{2}(1)=12.54;P<0.001$). ChatGPT 3.5 exhibited a smaller change, declining from 50.7% to 34.7%($\delta=16.0\%;\chi^{2}(1)=2.00;P=0.157$).
Figure 6: Accuracy of the LLMs when answering the CRT2 problems (A), and the proportion of incorrect answer types in modified CRT2 problems (B).
Upon scrutinizing all erroneous answers to the modified problems, as shown in Fig. 6B, it was observed that 52.8% of errors employed the original CRT2 methodology to address the modified problems (Maintain the original solution steps). This indicates that, due to the similarity in wording between the modified problem and the original problem, LLMs overlooked the differences in their mathematical principles and consequently chose the same solution steps as those for the original problem, which was incorrect.
@@ 3.3 For CRT3: Transforming Exponential Growth Problems into Linear Growth Problems
Fig. 7 presents the first problem in the CRT3 dataset as an illustrative example. In the original problem, the number of viruses doubles each day, resulting in exponential growth of the total number. In contrast, the modified problem features a constant daily increase in the number of viruses, which leads to linear growth. We calculated the average similarity using the same function, yielding a result of 88.89%. To further elucidate the problem-solving strategies of the LLM, we added a CoT prompt for both the original and the modified problem.
@@
The experimental results in Fig. 8A demonstrate that Gemini-1.5 achieved the highest accuracy of 100% for the original CRT3 problems. In contrast, ChatGPT 3.5 recorded the lowest accuracy at 54.0%, with an average accuracy of 86.8%. Following the modification of the problems, the average accuracy decreased significantly to 12.5% ($\delta=64.3\%;\chi^{2}(1)=272.83;P<0.001$). The accuracy rate of Gemini-1.5 decreased to 20.0%($\delta=80.0\%;\chi^{2}(1)=63.38;P<0.001$), while Claude3’s accuracy was merely 1.3% ($\delta=92.0\%;\chi^{2}(1)=81.23;P<0.001$).
Figure 8: Accuracy of the LLMs when answering the CRT3 problems (A), and the proportion of incorrect answer types in modified CRT3 problems (B).
The main reason for the decline in the average accuracy is that these models continued to employ the problem-solving approach of the original problem when confronted with the modified problem. The proportions of Claude3, ChatGPT 3.5, GPT-4, Gemini-1.5, and GLM-4 that identified the modified problem as one of exponential growth were 81.7%, 77.8%, 85.7%, 72.5%, and 71.5%, respectively, although it should have been identified as a linear growth problem. Among these, the proportion of problems that were solved entirely using the original method—yielding the same answer as the original problem—constituted 37.8%, 31.8%, 82.7%, 55.7%, 52.2% and the average is 51.4% (as depicted in Fig. 8B). This result indicates that more than half (51.4%) of the errors were due to maintaining the original solution steps.
Result Analysis of Experiment II: The modifications to CRT1, CRT2, and CRT3 altered the mathematical principles underlying these problems with a effort to preserve the similarity of their statements. The experimental findings reveal that LLMs continue to rely on their inherent problem-solving methodologies, even when provided with CoT prompts as guidance. This observation further supports our hypothesis that LLMs are capable of addressing mathematical problems, likely due to the inclusion of analogous problems in their training datasets. Consequently, LLMs select solutions based on the superficial similarity of problem descriptions, rather than demonstrating a deep understanding of the underlying mathematical principles.